OpenStudy (anonymous):

Question: 6 of 36: What is 8 5 ÷ 40 15 ?

OpenStudy (anonymous):

40/5=8

OpenStudy (anonymous):

this is the problem 8/5 % 40/15

OpenStudy (anonymous):

can some one help

OpenStudy (anonymous):

\[\frac{ 8 }{ 5 }*\frac{ 40 }{ 15 }*\frac{ 1 }{ 100 }=\frac{ 64 }{ 1500 }=0.042667\]